OpenStudy (anonymous):

When you simplify each side of the equation 4(x + 3) = 5 – (x + 1), you will get the equivalent equation 4x + 12 = -x + 4. TRUE OR FALSE ?

OpenStudy (lyrae):

4(x + 3) = 5 – (x + 1) 4x + 4*3 = 5 - x - 1 4x + 12 = -x + 4 True